OpenStudy (anonymous):

There are 48 cars and toy bicycles in a box. If there are a total of 128 wheels, how many toy bicycles are in the box?

OpenStudy (campbell_st):

solve get the 2 equations total number \[c + b = 48\] wheels \[4c + 2b = 128\] take the 1st equation and make b the subject \[b = 48 - c\] and substitute it into the 2nd equation. Now you can solve for c.

OpenStudy (anonymous):

thank you so much @campbell_st! Using your equation, there are 16 cars and 32 bicycles! Cheers!